OpenStudy (anonymous):

How many seven letter words can be formed without repetition from the letters of the word INCLUDE. That have the letters N and D separated by exactly two letters

hero (hero):

I would be able to do this, but that last part threw me off

OpenStudy (anonymous):

lol i got the first part already its just 7!

hero (hero):

I don't even get the last part. This is a dumb question

hero (hero):

N and D are separated by 3 letters, not two

OpenStudy (anonymous):

tell that to my text book lol

OpenStudy (anonymous):

thats isn't relevant because all the letters can rearranged there not in order

hero (hero):

Oh okay. Well I think you may have to do this by hand to figure it out

hero (hero):

So, I guess the question is, from the letters INCLUDE, how many ways can n and d be separated by 2 letters?

OpenStudy (anonymous):

yes I'm sorry i didn't write it correctly

OpenStudy (anonymous):

does it assume that letters cannot be repeated?

hero (hero):

Yes, no reps

hero (hero):

But I'm not so confident I can do it

OpenStudy (anonymous):

well i have the answer i just don't know how to do it

OpenStudy (anonymous):

do you reckon you could have a go at it?

hero (hero):

What's the answer?

OpenStudy (anonymous):

960

hero (hero):

Hmm, okay

hero (hero):

Yeah, clearly doing it by hand is not an option

OpenStudy (anonymous):

lol omg i hate probability

hero (hero):

I can't help you with this. It is beyond me

OpenStudy (anonymous):

i see well thanks for your effort

hero (hero):

I might post an explanation later.....we'll see

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

N _ _ D _ _ _N _ _ D _ _ _ N _ _ D

OpenStudy (anonymous):

and of course 3 more with the N and D switched.

OpenStudy (anonymous):

so how many if it looks like N _ _ D _ _ ?

OpenStudy (anonymous):

you have 4 slots to fill in so 4!

OpenStudy (anonymous):

and likewise for _ N _ _ D _ and for _ _ N _ _ D

OpenStudy (anonymous):

each of these have 4! ways

OpenStudy (anonymous):

so unless i am totally off we have \[6\times 4!\] ways of doing it, by counting. does that look reasonable?

OpenStudy (anonymous):

i get 144

OpenStudy (anonymous):

unfortunately its 960 lol

OpenStudy (anonymous):

damn hold on let me think!

OpenStudy (anonymous):

ok lol:)

OpenStudy (anonymous):

oh because i can't count!

OpenStudy (anonymous):

there are 7 letters not 6!

OpenStudy (anonymous):

i am a moron yikes. it is not \[6\times 4!\] it is \[8\times 5!\]

OpenStudy (anonymous):

if i could count correctly it would be clear so lets start again. we have 7 letter. could look like N _ _ D _ _ _ _ N _ _ D _ _ _ _ N _ _ D _ _ _ _ N _ _ D

OpenStudy (anonymous):

or the same with the N and D switched.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so 8 possibilities all together

OpenStudy (anonymous):

each of which has 5! arrangements because you are filling 5 slots

OpenStudy (anonymous):

so the answer is \[8\times 5!\]

OpenStudy (anonymous):

now i hope this is what you have

OpenStudy (anonymous):

thank you so much ur awesome:)

OpenStudy (anonymous):

960 right?

OpenStudy (anonymous):

yep thats correct :)

OpenStudy (anonymous):

yes i scrolled up so it is right

OpenStudy (anonymous):

we could have done it quickly if i knew how to count to 7 instead of 6 don't you hate combinatorics?

OpenStudy (anonymous):

yes i do lol i hate probability in general would much rather do calculus:)

hero (hero):

Yeah, I figured it out too.